Tohban Report 2018-03-21
Tohban Reports | |
---|---|
Start Date: | 14 March 2018 |
End Date: | 21 March 2018 |
Tohban: | Jim McTiernan |
Tohban email: | jimm@ssl.berkeley.edu |
Next Tohban: | TBD |
List all reports |
Solar Activity
Quiet...
No RHESSI flares to count. Maybe an A5 flare for GOES on 20-mar-2018.
Memory Management
Decimation was active/vigorous for the full time period, SSR fill was up to 83% on 2018-03-15, and the pointer was jumped to free all of the memory. This resulted in a Vc3 data gap from 7-mar-2018 02:56 until 15-mar-18:45. Currently the SSR fill is approximately 50%.
Spacecraft Status
Detectors 3 and 8 are recording events. The cold tips are 136.4 and 134.6 K as of 19-Mar-2018. The cold plates, 1 and 2 are 159.6 and 158.2 K. Cryocooler power is 54.9W, cryocooler efficiencies are 0.0339 and 0.0343. See:
http://sprg.ssl.berkeley.edu/~ayshih/soh/latest_temperatures.pdf
for a summary plot of recent temperatures, efficiencies and accelerometer values.
Data Gaps
DATA GAPS FOR 14-mar-2018 TO 22-mar-2018, note the the large gaps for 21-march-2018 are likely to be filled.
TIME RANGE: 2018-03-14T00:00:00.000 -- 2018-03-22T00:00:00.000 GAPS IN APP_ID = 1 (VC1-SOH) WITH PACKET RATE LT 30 N_GAPS 1 GAP START TIME GAP END TIME GAP (SEC) 2018-03-21T15:50:00.000 -- 2018-03-22T00:00:00.000 29400.000 GAPS IN APP_ID = 154 (VC1-PMTRAS) WITH PACKET RATE LT 4 N_GAPS 5 GAP START TIME GAP END TIME GAP (SEC) 2018-03-18T05:10:00.000 -- 2018-03-18T10:50:00.000 20400.000 2018-03-19T11:05:00.000 -- 2018-03-19T12:00:00.000 3300.0000 2018-03-20T10:35:00.000 -- 2018-03-20T11:30:00.000 3300.0000 2018-03-21T12:00:00.000 -- 2018-03-21T12:40:00.000 2400.0000 2018-03-21T15:50:00.000 -- 2018-03-22T00:00:00.000 29400.000 GAPS IN APP_ID = 102 (VC3-MONITOR RATES) WITH PACKET RATE LT 30 N_GAPS 6 GAP START TIME GAP END TIME GAP (SEC) 2018-03-14T00:00:00.000 -- 2018-03-15T18:45:00.000 153900.00 2018-03-15T23:55:00.000 -- 2018-03-16T01:10:00.000 4500.0000 2018-03-16T07:20:00.000 -- 2018-03-16T10:45:00.000 12300.000 2018-03-16T18:25:00.000 -- 2018-03-16T18:30:00.000 300.00000 2018-03-17T22:45:00.000 -- 2018-03-18T02:25:00.000 13200.000 2018-03-18T17:05:00.000 -- 2018-03-22T00:00:00.000 284100.00
Detector issues
Detectors 3 and 8 are powered on, and are recording events. No threshold changes or HV changes this past week.
Spacecraft Management
Decimation | Active/Vigorous |
HLAT Decimation | Rear decimation |
Night time data (fronts) | plus/minus 4 minutes from daylight |
Night time data (rears) | plus/minus 4 minutes from daylight |
Require extra passes? | No |
Requirement for moving pointer? | See above |
Attenuator operation | Nominal |
Ref. deadtime for Att.thresholds | 22.x% |
Detector problems? | See above |